# VisionUnite: A Vision-Language Foundation Model for Ophthalmology Enhanced with Clinical Knowledge

### v1

@@ Methods
We employ three distinct baseline methods for comparative analysis with VisionUnite, GPT-4V [20], Gemini Pro [21], and manual interpretation from the junior ophthalmologist. All methods are evaluated on a test set comprising 180 images. Notably, the manual interpretation involves diagnostic opinions and bases generated by junior ophthalmologists. Subsequently, the outcomes from the three models and junior ophthalmologists undergo an assessment by senior ophthalmologists to evaluate the accuracy and relevance of the generated text.
Implementation details
VisionUnite is implemented using Python 3.8.18 and Pytorch ’2.0.1+cu117’. The fine-tuning of VisionUnite is run on 8-card NVIDIA A100 80G and 128 Intel(R) Xeon(R) Platinum 8369B CPU @ 2.90GHz. The batchsize we set is 4 per GPU in each mini-barch and the accumulated gradient iterations is 1, so the VisionUnite is trained with a batch size of 32. We utilize the AdamW optimizer with base lr=0.001 and the betas in AdamW is set as (0.9, 0.95). The absolute lr, which equals $baselr*batchsize/256=1.25E-4$. The lower lr bound for cyclic schedulers is 0 and the weight decay is 0.02. The number of training epochs is 10 in the pre-training stage. The number of training epochs is 30 in the fine-tuning stage and the first epochs are set for warming up. The number of workers we set is 10. The max token length is 512 to ensure that all the text can be included. We adopt the DistributedDataParallel (DDP) as the data parallel mechanism during training.
Evaluation and statistical analysis
The diagnoses and explanations provided by VisionUnite, Gemini Pro, GPT-4V, and the junior ophthalmologist are scrutinized by senior ophthalmologists to determine the correctness of different outcomes. Evaluation metrics such as diagnostic accuracy and diagnostic relevance are employed for statistical analysis. As for diagnostic relevance, ophthalmologists rank the correlation between the predictions of the four groups and the actual situation, assigning a relevance score from 1 (least relevant) to 4 (most relevant). The average diagnostic relevance is calculated statistically to depict the diagnostic performance of different methods. It is worth noting that all three rounds of dialogue response are utilized for prediction analysis. We also analyze the misdiagnosis situation of the models and classified the errors into minor errors and major errors according to the severity. We calculate the 95% confidence interval for each result and the p-value between it and our proposed VisionUnite. Furthermore, we design four scenarios of experiments to probe into the diagnosis of challenging cases and the error correction capabilities of models. In the first scenario, specific disease and condition samples are selected to investigate the generalization ability of models for difficult scenarios. In the second scenario, we investigate the diagnostic ability of models for fundus images with multiple types of diseases or conditions. In the third scenario, the misdiagnosis situation is explored for healthy fundus images. In the fourth scenario, we analyze the changes in diagnostic results and overall response quality of different models in the multiple rounds of dialogue to explore their correction abilities with patients. In addition, we analyze the diagnostic errors. The diagnostic errors are divided into missed errors and incorrect errors and each error is classified into two error levels: minor and major. Finally, we also investigate the performance of VisionUnite assisting junior ophthalmologists and the consistent interpretation of visual and language Features in different models. These experiments serve to further verify the generalization ability and robustness of VisionUnite.
Data availability
The pretrain natural image-text pairs are from the MS COCO dataset [12], which is available at https://cocodataset.org/#download. The pretrain biomedical image-text pairs are from the Retina Image Bank dataset [15] available at https://imagebank.asrs.org/, and the PMC-OA dataset [13, 14], which is available at https://huggingface.co/datasets/axiong/pmc_oa.
The MMFundus dataset can be accessed at https://github.com/HUANGLIZI/MMFundus, and its image part comprises the following datasets: 1. HRF dataset [33], 2. INSPIRE-AVR dataset [34], 3. IOSTAR dataset [35], 4. RITE dataset [36], 5. G1020 dataset [37], 6. GAMMA dataset [38], 7. ORIGA dataset [39], 8. REFUGE dataset [19], 9. ODIR dataset [16], 10. PALM dataset [40], 11. RFMiD dataset [41], 12. RFMiDv2 dataset [42], 13. APTOS dataset [17], 14. DeepDRiD dataset [18], 15. EyePACS dataset [43], 16. IDRID dataset [44], 17. ADAM datase [45], 18. ACRIMA dataset [46], 19. MESSIDOR-2 dataset [47], 20. JSIEC dataset [48], 21. DeepEyeNet dataset [49], 22. LAG dataset [50], 23. PARAGUAY dataset [51], 24. PAPILA dataset [52], 25. STARE dataset [53], 26. FIVES dataset [54], 27. FUND dataset [55], 28. E-ophta dataset [56], 29. BRSET dataset [57], 30. MuReD dataset [58], 31. OIA-DDR dataset [59], 32. SUSTech-SYSU dataset [60], 33. Cataract dataset [61], 34. DGOCF dataset [62], 35. BoVW dataset [63], 36. HarvardGlaucoma dataset [64], 37. RIM-ONE dataset [65], 38. CHAKSU dataset [66], 39. DiaRetDB dataset [67], 40. LSD dataset [68], 41. GNG dataset [69], 42. AOD dataset [70], 43. DHRF dataset [71], 44. VietAI dataset [72], 45. ToxoFundus dataset [73], 46. Papilledema dataset [74], 47. BEH dataset [75], 48. WilliamHoyt dataset [76], 49. ROI dataset [77], 50. ROD dataset [78], 51. BiDR dataset [79], 52. AIROGS dataset [80], 53. Private dataset. The private dataset is from a hospital in China mainland, which includes two major types of diseases "myopia" and "glaucoma", as well as clinical features such as the cup-to-disc ratio and intraocular pressure. The characteristics of patients in the private dataset are in the supplementary material.
Code availability
The code of VisionUnite for the training and evaluation is available on GitHub at https://github.com/HUANGLIZI/VisionUnite.
Acknowledgements
This research is supported by Shanghai Artificial Intelligence Laboratory. This work is partially supported by the National Key R&D Program of China(NO. 2022ZD0160100), partially supported by the National Natural Science Foundation of China (62272450), and partially supported by the China Postdoctoral Science Foundation (Grant No.2022M721743). This work was done during Zihan Li’s internship at Shanghai Artificial Intelligence Laboratory.
Author contribution statement
Z.L., D.S., and Y.Q. conceived and designed the study. Z.L., D.S. Z.Y., D.W., F.L., and X.Z. acquired the data. Z.L., Z.Y., and D.W. did the statistical analyses. Z.L., D.S. developed, trained, and applied the foundation model. Z.Y., D.W., F.L., and X.Z. verified the underlying raw data; Z.L. and D.S. prepared the first draft of the manuscript; X.Z., P.K., and Y.Q. conceived and oversaw the project, interpreted data, reviewed the manuscript, and acted as guarantor. All authors contributed to manuscript preparation and modification. All authors analyzed and interpreted the data. All authors provided suggestions on the framework.
Figure 1: (a) Previous vision models could only diagnose specific diseases as positive or negative, lacking the ability to provide clinical explanations or interact with patients. However, our proposed VisionUnite changes this approach. It can predict a wide range of diseases and allows real-time conversations with patients, incorporating their feedback. Additionally, VisionUnite offers clear clinical explanations in its output, making it more understandable and useful. (b) The label distribution of the proposed MMFundus dataset, which includes eight main categories excluding the "Others" class. (c) VisionUnite is built with a transformer-based vision encoder and a specialized vision adapter designed for classifying six different signs including Vascular, Macular, FBC (Fundus Boundary Color), OCD (Optical Cup Disc), FHE (Fundus Hemorrhages Examination), and Other. It includes a vision projector to align visual embeddings with text tokens. (d) The illustration of image-text contrastive learning (CLIP Loss). (e) The illustration of classification supervised learning (CLS Loss). (f) The illustration of text-generation supervised learning (LLM Loss).

### v2

@@ IV-A2 Baseline methods
We employ three distinct baseline methods for comparative analysis with VisionUnite, GPT-4V [9], Gemini Pro [8], and manual interpretation from the junior ophthalmologist. All methods are evaluated on a test set comprising 180 images. The manual interpretation involves diagnostic opinions and bases generated by junior ophthalmologists. The outcomes from three models and junior ophthalmologists undergo an assessment by senior ophthalmologists to evaluate the accuracy and relevance of the generated text.
IV-A3 Evaluation and statistical analysis
We evaluate model efficacy using a multi-round VQA dataset from the MMFundus dataset, consisting of 180 samples and 540 question rounds. This dataset includes a spectrum of fundus conditions, from healthy to thirteen different ophthalmic diseases. Our evaluation benchmarks model performance against open-set clinical diagnostics based on two key criteria: diagnostic accuracy and relevance. Diagnostic accuracy measures the precision of disease identification and categorization. However, since accuracy alone may not fully capture a model’s performance, we introduce diagnostic relevance to assess the quality of model responses and distinguish between different models. Our testing setup involves presenting a fundus image and three related questions to a junior ophthalmologist with one year of experience and eliciting responses from three advanced vision-language models: GPT-4V [9] (gpt-4-1106-vision-preview), Gemini Pro [8] (gemini-pro-vision), and our proposed VisionUnite. These responses are then assessed by three senior ophthalmologists with over ten years of experience, focusing on both diagnostic accuracy and relevance. We statistically calculate the average diagnostic relevance to compare the different models. Additionally, we analyze model misdiagnoses, categorizing errors as minor or major based on their severity. We also compute the 95% confidence intervals and p-values for each model’s results, particularly comparing them to our VisionUnite model, ensuring that responses are integrated into coherent paragraphs without additional AI-generated content for fair evaluation. We also design a multiple-choice evaluation experiment to evaluate the diagnostic performance across various models in a close-set setting. We conduct experiments on 2233 cases and calculate the multiple-choice accuracy of ten diseases. The details are as follows:
1. Diagnostic accuracy: We follow the clinical evaluation standards. 1. In cases of extra answers, we consider them to be incorrect answers. 2. In case of just missing unnecessary diagnostic information, we still consider the responses to be correct. 3. In our evaluation, the answer must only include all diagnosable diseases to be considered correct. We use the Wilson method [76] to estimate 95% confidence interval of diagnostic accuracy and calculate the p-value with the two-sided T-test and the above Wilson estimation.
2. Diagnostic relevance: Senior ophthalmologists rank the response sets based on their alignment with the accurate diagnosis, scoring the most consistent response as four and the least as one. The diagnostic relevance is designed to refine performance evaluation by taking into account factors that mere diagnostic accuracy might miss. For example, a response could be accurate in diagnostic prediction yet fail to adhere strictly to diagnostic criteria, which will diminish its overall diagnostic relevance. The bootstrap [77] method is used to estimate the 95% confidence interval of diagnostic relevance and the two-sided T-test is applied to calculate its p-value.
3. Sentence-BERT (SBERT) Similarity: Our evaluation also employs Sentence-BERT (SBERT) Similarity [78] as the quantitative metric for assessing semantic correspondence between generated responses and ground truth. We report the results of SBERT Similarity across 540 multi-round dialogue instances.
4. Multiple-choice accuracy: We provide four distinct options from all disease labels in the MMFundus dataset that models should select for the most likely diagnostic outcome based on the provided fundus image. The other three options are randomly selected except the correct options.
IV-B Comparison of Diagnosis between Ophthalmologist and Large Vision-Language Models
In our study, we conduct a comprehensive assessment of the diagnostic capabilities exhibited by the junior ophthalmologist compared with models, which include Gemini Pro, GPT-4V, and our proposed VisionUnite. The evaluation is performed on the designated test dataset. Additionally, we explore an examination of the diagnostic relevance between the responses provided by the junior ophthalmologist and those generated by models. As shown in Table II, the overall diagnostic accuracy of VisionUnite is over 45% and 28% higher than Gemini Pro and GPT-4V respectively, and the corresponding p-value is less than 0.001. The diagnostic relevance of VisionUnite is also over 62% and 25% higher than Gemini Pro and GPT-4V respectively with the p-value less than 0.001. Compared to the junior ophthalmologist, the overall accuracy of VisionUnite is approximately 4.5% higher with the p-value being 0.0876, and the diagnostic relevance is 0.75% higher (from 2.915 to 2.937). In each round of result evaluation, we find that the performance of VisionUnite in the first round of diagnostic Q&A is higher than that of junior ophthalmologist, with an accuracy rate of approximately 7.2% and a diagnostic relevance of 2.6%. The result indicates that VisionUnite has stronger analytical and reasoning abilities for fundus images, thus achieving superior performance in the first round of diagnosis. In the second and third rounds of diagnostic Q&A, the performance of VisionUnite is comparable to that of a junior ophthalmologist and far superior to GPT-4V and Gemini Pro. It also indicates that VisionUnite has the same problem-solving and interpretation abilities as doctors.